OpenStudy (anonymous):

if b > a > 1, then ??? please explain!

OpenStudy (anonymous):

explain what?

OpenStudy (anonymous):

oh wait- here are the answer choices

OpenStudy (stormfire1):

then b > 1. It's a logic question. If a > 1 and b is > a, then b must be > 1.

OpenStudy (stormfire1):

what are the choices? There can be other answers....

OpenStudy (anonymous):

In english: If a is between b and 1 then a> b and a < 1

OpenStudy (anonymous):

a. 1/a < 1/b b. a^2 < b^2 c. n < 3 d. a + 2b > 3b e. b^2 < ab

OpenStudy (anonymous):

anyone know?????

OpenStudy (anonymous):

b

OpenStudy (anonymous):

I like the first option.

OpenStudy (stormfire1):

b

OpenStudy (anonymous):

why is it b??

OpenStudy (anonymous):

b>a or a<b taking square on both side will make b option

OpenStudy (anonymous):

b > a and a > 1 b > a and a< b

OpenStudy (anonymous):

\[\text{let } a = \frac{1}{2}, b=\frac{1}{4}\] \[\implies b < a < 1\] \[b^2 = \frac{1}{16}, a^2 = \frac{1}{4}\] \[a^2 \not< b^2\]

OpenStudy (stormfire1):

Actually, on second thought, it's a....because a or b can be negative.

OpenStudy (anonymous):

Indeed

OpenStudy (anonymous):

so it is A?

OpenStudy (stormfire1):

Most definitely.

OpenStudy (stormfire1):

Of course, C works too...if n is in fact < 3 :P

OpenStudy (anonymous):

it can only be one answer though

OpenStudy (stormfire1):

C was a *joke* :) It's A

OpenStudy (anonymous):

\[a= \frac{1}{2}, b= -5\]\[\implies b < a < 1\]\[\frac{1}{a} = 2, \quad\frac{1}{b} = -\frac{1}{5}\]\[\frac{1}{a} > \frac{1}{b}\].. Hrm it's not a.

OpenStudy (anonymous):

at polpak read the question carefully it is b>a>1 not b<a<1

OpenStudy (anonymous):

Oh, botheration. Why would you order it that way?

OpenStudy (anonymous):

so option b is correct

OpenStudy (anonymous):

does anyone have a final answer thats RIGHT?

OpenStudy (anonymous):

im going with b

OpenStudy (anonymous):

b is correct. I misread the initial inequality.

OpenStudy (anonymous):

Sorry for the confusion.

OpenStudy (anonymous):

its ok thanks for the help

OpenStudy (stormfire1):

If a = -3 and b = -2, then a^2 = 9 and b^2=4....so B cannot be right.

OpenStudy (anonymous):

ohhh how -2 is grater than 1

OpenStudy (stormfire1):

ah crap...I forgot about that. Ok, B it is

OpenStudy (anonymous):

it is given in question

OpenStudy (stormfire1):

Never change your first answer...that's what I should have done :)